a card is taken at random from a pack of cards and replaced. A second card is drawn. The probability of a diamond being drawn is one quarter. What is the probability that at least one card is a diamond?
Answer by stanbon(75887)

a card is taken at random from a pack of cards and replaced. A second card is drawn. The probability of a diamond being drawn is one quarter. What is the probability that at least one card is a diamond?
---------------------------------
P(one is diamond) = 1 - P(none is diamond)
---
= 1 - (3/4)^2
---
= 1 - 9/16
---
= 7/16
============
cheers,
Stan H.